Solve for x when 3^x=9x
Found 3 solutions by ultrakimidis1, Alan3354, ikleyn:
Answer by ultrakimidis1(8)

3^x=9
3^x=3³
Equating exponents
X=3

Answer by Alan3354(69443)

Solve for x when 3^x=9x
x = 3 by thinking about it.
==========
3^x=9x
An algebraic solution is not possible.
You can use graphs or iteration methods.
-----
If it were
3^x = 10x, no solution is obvious, and one of the approximation methods is necessary.


Answer by ikleyn(52788)

.
The solution is x= 3.